# Efficient generation of divergent and collimated hot electrons via a novel multi-beam two-plasmon decay and stimulated Raman scattering mechanism

### v1

@@ IV.1 Effects of asymmetric incidence
In realistic experiments on large-scale laser facilities, despite the symmetric configuration of the beams, the orientation of density contours evolves in space and time during the target surface ablation. As a result, it is common for laser beams to deviate from symmetry relative to the normal of the density contours near $n_{c}/4$. Therefore, hot electron generation under asymmetric incidence is crucial for accurately assessing $f_{hot}$ in experimental evaluations. In such scenarios, the matching ellipses [Figure (4a)] for both TPD and SRS mechanisms continue to define the common wave modes responsible for the TSI mechanism. The cumulative wave amplification at various densities can still be scaled by the total gain, $G_{tot}$, as derived from equations (7) to (10).
@@
IV.2 Experimental observability of the TSI mechanism
Since the scattered light in the TSI mechanism is essentially enhanced SRS scattered light, it should be detectable in experiments. For the simulation (i) [table 1] with symmetric incident angle $\theta_{in}=52^{\circ}$, if we assume $n_{e}$ only varies along $x-$ direction in space, the scattered light with a frequency of $\omega\sim 0.512\omega_{0}$ and wave vector component $k_{y}\sim 0.212\omega_{0}/c$ will ultimately have a scattering angle of only $\sim 24^{\circ}$ relative to the negative x-direction after exiting into the vacuum. This scattered light has the same frequency as the collective Thomson scattering of the incident light on the common EPW mode. Once diagnosed, it should mix with the double-band structure of the TPD $\omega_{0}/2$ spectra, and strengthen the blue-shift component. This qualitatively agrees with the observations of significantly stronger blue-shift components compared to red-shift components off the target normal direction in previous OMEGA experimentsSeka et al. (2014).
@@ IV.3 The actual incident angles in the vacuum
IV.4 The connection between picosecond PIC simulation results to nanosecond hydrodynamic evolution
A key question is the extent to which our PIC simulation results, obtained on a picosecond time scale, can inform the realistic implosion process occurring over several nanoseconds. The significant disparity between these time scales naturally raises concerns about the applicability of our results to real-world experiments. To address this, it is important to note that as long as a quasi-steady state is achieved during the short-time-scale process, quantities from this state can reasonably be coupled to the long-time-scale process with slowly evolving hydrodynamic parameters. This is because the long-time-scale process is not influenced by transient growth or decay occurring before the system reaches its quasi-steady state. Following this principle, previous studiesFollett et al. (2017) have successfully predicted the time evolution of hot electron energy in nanosecond-scale experiments by using results from many hybrid fluid-particle simulations (each runs for 20 ps) that model hot electron generation under various laser-plasma conditions throughout the entire laser pulse and focal spot. Another successful example is the development of iterative methods that couple 15 ps-PIC and 2 ns-hydro simulations to assess the impact of TPD-induced strong pump depletion due to anomalous laser absorption and hot electron generation on hydrodynamic evolutionLi, Hu, and Ren (2017).
In the context of our simulations, since the results clearly exhibit quasi-steady states [figure (2)], we expect our findings to reliably reflect the hot electron characteristics for the corresponding physical parameters and beam geometry of the laser pulses and plasmas.
@@ IV.4 The connection between picosecond PIC simulation results to nanosecond hydrodynamic evolution
In this study, we investigate the angular variations of hot electrons generated by multiple laser beams with varying incident angles. We identify a novel multi-beam collaborative process involving two-plasmon decay (TPD) and stimulated Raman scattering (SRS), termed TSI, which produces hot electrons with distinct divergent and collimated components.
We demonstrate that the key factor influencing the acceleration of these components is the common Langmuir wave associated with TSI. By properly modeling the common wave gain, we obtain scaling relations between the gains and the energy fractions of hot electrons at different angles. Our results show that the energy of divergent hot electrons is more sensitive to variations in the common wave gain. Furthermore, when the density gradient deviates from the bisector of the laser beams— a common occurrence in large-scale laser facilities—our study reveals that asymmetric beam incidence results in unequal common wave gains, leading to asymmetric angular distributions of hot electrons. The stronger peak is observed in the direction opposite to the laser beam with the larger incident angle. These findings highlight potential preheating concerns due to divergent and collimated hot electrons in laser-plasma configurations with high gains in future DCI experiments.
Finally, we observe that the scattered light associated with the TSI mechanism qualitatively agrees with previous experimental findings. This scattered light shares the frequency of collective Thomson scattered light from the common wave. Although the onsite scattering angle is large (approximately $95^{\circ}$), the direction of the scattered light shifts towards the density gradient when propagating towards vacuum, with final angle of about $24^{\circ}$ in our simulations for 1D plasma density variations. This scattered light merges with the double-band structure of the TPD $\omega_{0}/2$ spectrum, enhancing the blue-shift component, consistent with prior OMEGA experimentsSeka et al. (2014) that observed significantly stronger blue-shift than red-shift signals when viewed off the target normal.
@@ Acknowledgements.
This research was supported by the National Natural Science Foundation of China (NSFC) (Grant Nos. 12275269, 12175229, 12275032), by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Nos. XDA25010200 and XDA25050400). The numerical simulations in this paper were conducted on Hefei advanced computing center.

### v2

@@ IV.1 Effects of asymmetric incidence
In realistic experiments on large-scale laser facilities, despite the symmetric configuration of the beams, the orientation of density contours evolves in space and time during the target surface ablation. As a result, it is common for laser beams to deviate from symmetry relative to the normal of the density contours near $n_{c}/4$. Therefore, hot electron generation under asymmetric incidence is crucial for accurately assessing $f_{hot}$ in experimental evaluations. In such scenarios, the matching ellipses [Figure (4a)] for both TPD and SRS mechanisms continue to define the common wave modes responsible for the STS mechanism. The cumulative wave amplification at various densities can still be scaled by the total gain, $G_{tot}$, as derived from equations (7) to (10).
@@
IV.2 Experimental observability of the STS mechanism
Since the scattered light in the STS mechanism is essentially enhanced SRS scattered light, it should be detectable in experiments. For the simulation (i) [table 1] with symmetric incident angle $\theta_{in}=52^{\circ}$, if we assume $n_{e}$ only varies along $x-$ direction in space, the scattered light with a frequency of $\omega\sim 0.512\omega_{0}$ and wave vector component $k_{y}\sim 0.212\omega_{0}/c$ will ultimately have a scattering angle of only $\sim 24^{\circ}$ relative to the negative x-direction after exiting into the vacuum. This scattered light has the same frequency as the collective Thomson scattering of the incident light on the common EPW mode. Once diagnosed, it should mix with the double-band structure of the TPD $\omega_{0}/2$ spectra, and strengthen the blue-shift component. This qualitatively agrees with the observations of significantly stronger blue-shift components compared to red-shift components off the target normal direction in previous OMEGA experimentsSeka et al. (2014).
@@ IV.3 The actual incident angles in the vacuum
In this study, we investigate the angular variations of hot electrons generated by multiple laser beams with varying incident angles. We identify a novel multi-beam collaborative process involving two-plasmon decay (TPD) and stimulated Raman scattering (SRS), termed STS, which produces hot electrons with distinct divergent and collimated components.
We demonstrate that the key factor influencing the acceleration of these components is the common Langmuir wave associated with STS. By properly modeling the common wave gain, we obtain scaling relations between the gains and the energy fractions of hot electrons at different angles. Our results show that the energy of divergent hot electrons is more sensitive to variations in the common wave gain. Furthermore, when the density gradient deviates from the bisector of the laser beams— a common occurrence in large-scale laser facilities—our study reveals that asymmetric beam incidence results in unequal common wave gains, leading to asymmetric angular distributions of hot electrons. The stronger peak is observed in the direction opposite to the laser beam with the larger incident angle. These findings highlight potential preheating concerns due to divergent and collimated hot electrons in laser-plasma configurations with high gains in future DCI experiments.
Finally, we observe that the scattered light associated with the STS mechanism qualitatively agrees with previous experimental findings. This scattered light shares the frequency of collective Thomson scattered light from the common wave. Although the onsite scattering angle is large (approximately $95^{\circ}$), the direction of the scattered light shifts towards the density gradient when propagating towards vacuum, with final angle of about $24^{\circ}$ in our simulations for 1D plasma density variations. This scattered light merges with the double-band structure of the TPD $\omega_{0}/2$ spectrum, enhancing the blue-shift component, consistent with prior OMEGA experimentsSeka et al. (2014) that observed significantly stronger blue-shift than red-shift signals when viewed off the target normal.
@@ Acknowledgements.
This research was supported by the National Natural Science Foundation of China (NSFC) (Grant Nos. 12275269, 12175229, 12375243, 12275032, 12388101), by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Nos. XDA25010200 and XDA25050400). The numerical simulations in this paper were conducted on Hefei advanced computing center.